OpenStudy (anonymous):

Integrate by trig sub. \[\int\limits_{}^{}dx/x^{2}\sqrt{4-x^2}\] let x=2sintheta

OpenStudy (anonymous):

\[\int\limits_{}^{}dx/x^{2}\sqrt{4-x^2}\]

OpenStudy (anonymous):

Do you know how to do a \(u\) sub? Trig sub isn't really much different.

OpenStudy (anonymous):

Find dx

OpenStudy (anonymous):

Yes, I know how to do it all. I do all the sub, and after doing it all, i can get it down to .25\[.25\int\limits_{}^{}1/\sin^2\theta\]

OpenStudy (anonymous):

now i have to go back into terms of x, and thats where im having trouble

OpenStudy (anonymous):

You have to do the integral first.

OpenStudy (anonymous):

What is \(\frac{d(\tan(\theta))}{d\theta}\)? What is \(\frac{d(\cot(\theta))}{d\theta}\)?

OpenStudy (anonymous):

1/sin^2theta = csc^2theta. integral of csc^2theta = -1cot(theta)

OpenStudy (anonymous):

Okay, so before undoing the trig sub, what is the anti-derivative in terms of theta?

OpenStudy (anonymous):

\[\int\limits_{}^{}dtheta=\theta\]

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Isn't it \(-0.25\cot(\theta) +C\)

OpenStudy (anonymous):

yes, thats right

OpenStudy (anonymous):

\[\int\limits \frac{dx}{x^2\sqrt{4-x^2}}= \frac{1}{4}\frac{(x-2)(x+2)}{x \sqrt{4-x^2}}+c\]

OpenStudy (anonymous):

need to do it through trig sub mathsmind :<

OpenStudy (anonymous):

that's what i got, solving on piece of paper

OpenStudy (anonymous):

Now you need to remember that \(\cot = \cos/\sin\) and that \(\cos = \sqrt{1-\sin^2}\)

OpenStudy (anonymous):

oh ok that would make life easier like a piece of cherry on top of the cake

OpenStudy (anonymous):

answer in the back of the book is \[-\sqrt{4-x^2}/4x+c\]

OpenStudy (anonymous):

yes that is right and my answer is right too...

OpenStudy (anonymous):

So basically: \[ 2\cot(\theta) = \frac{\sqrt{4-(2\sin(\theta))^2}}{2\sin(\theta)} \]

OpenStudy (anonymous):

because you are asked to use trig sub

OpenStudy (anonymous):

wio i dont have 2cot though, i have -.25cot so far and im trying to turn that into what i just posted

OpenStudy (anonymous):

\(-0.25 = 2 \times -0.125\)

OpenStudy (anonymous):

\( -0.25\cot = -0.125 \cdot 2\cot\)

OpenStudy (anonymous):

Could somebody just solve the problem for me, with the final answer being in terms of x and i can see where im going wrong? i essentially have majority of the problem right, but im messing up somewhere in the end

OpenStudy (anonymous):

wait am solving this on a piece of paper be patient

OpenStudy (anonymous):

i will be.

OpenStudy (anonymous):

oh wow. i figured it out haha. i realized where i was making my mistake :<

OpenStudy (anonymous):

ok do you agree that the final answer would be -cos(theta)/sin(theta)

OpenStudy (anonymous):

all you have to do is sub things back in

OpenStudy (anonymous):

so in our assumptions x^2=4sin^(theta)

OpenStudy (anonymous):

also we said 4-4sin^(theta)=4cos^(theta)

OpenStudy (anonymous):

i was forgetting the basic identity that 1/tan=cotan. using that identity, i was able to figure out where i was making the mistake, thanks mathsmind and wio

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

all you have to do is move backwords sub x's back in you'll get the exact answer in the book, i used partial fraction to solve the above problem which will lead to the same answer after simplification

OpenStudy (anonymous):

\[ \begin{split} \int\frac{dx}{x^{2}\sqrt{4-x^2}} &= \int\frac{2\cos(\theta)d\theta}{4\sin^{2}(\theta)\sqrt{4-4\sin^{2}(\theta)}} \\ &= \int \frac{d\theta}{4\sin^2(\theta) } \\ &=\int \frac{1}{4}\csc^2(\theta)d\theta \\ &= -\frac{\cot(\theta)}{4}+C\\ &=-\frac{1}{8}\frac{2\cos(\theta)}{2\sin(\theta)} +C \\ &= -\frac{1}{8}\frac{\sqrt{4-4\sin^2(\theta)}}{2\sin(\theta)} \\ &= -\frac{\sqrt{4-x^2}}{8x} \end{split} \]

OpenStudy (anonymous):

\(+C\)